Easy trick of finding cubes 3 ,2 digit number ?????????????????????????????

Grade:8

2 Answers

Arun
4 years ago

Find the cube of  728 ( 728³= ? )

Step 1 : Split the number as like i.e  7 – 28 or 72 – 8. 

Here we take 72 – 8 . 

Step 2 : Divide the 1st part by 2nd part. i.e 72/8 = 9. Here we can take a = 8 , b= 9 , c =  = 512

Step 3 : Cube of 1st part c = 8 x 8  x 8 = 512. Here Take last digit in answer and remaining value 51 to be add to next step.

i.e Ans :  _ _ _ _ 2  and 51 is to be add in the next step.

Step 4 : 3 bc + 51= (3 x 9 x 512) +51=  13824 + 51 = 13875 . Here Take lost digit into the next digit in answer and remaining value 1387 to be add to next step.

i.e Ans :  _ _ _ 5 2  and 1387 is to be add in the next step.

Step 5 : 3 b² c +13875= (3 x 9 x 9 x 512) +1387= 124416 + 1387 = 125803 . Here Take lost digit into the next digit in answer and remaining value 12580 to be add to next step.

i.e Ans :  _ _ _3 5 2  and  12580 is to be add in the next step.

Step 6 : b³ c +12580= ( 9 x 9 x 9 x512) +12580= 373248 + 12580 = 385828. Here Take this value into the answer

i.e Ans :  385828352.

Yashraj
3 years ago
there are three ways just learn as it is easy 
by vedic maths 
or other that we do basically....................